OpenStudy (anonymous):

i need someone to give me an example or other information about simpson 's rule?

OpenStudy (anonymous):

It's a method of numerical integration that uses parabolas instead of straight lines like rectangular or trapezoidal approximation.

OpenStudy (anonymous):

can you give some example about simpson's rule

OpenStudy (anonymous):

It's just like any approximation method. Test how quickly it converges to the answer for \[\int\limits_{0}^{\pi} \sin(x)dx\] when compared to rectangular approximation and triangular. Compare the errors for equal number of divisions of the interval.

OpenStudy (anonymous):

what is the first step on how to compute the example you'd given?

OpenStudy (anonymous):

Do you know the fomula for the rule? I'd say look up some info on the intuition. The part about the derivation of the parabola is neat.